OpenStudy (anonymous):

Find the slope and the y-intercept 9- 1/5y=10x

jimthompson5910 (jim_thompson5910):

9- 1/5y=10x -1/5y=10x-9 y=-5*10x+(-5)*(-9) y=-50x+45 Equation is now in y=mx+b form where m=-50 and b=45. So the slope is -50 and the y-intercept is (0,45)